A marksman fires a rifle horizontally at a target. How far does the bullet drop in flight if the target is 150 m away and the bullet has a muzzle velocity of 500 m/s?
1
Expert's answer
2020-11-16T07:50:58-0500

Let's first find the flight time of the bullet from the kinematic equation:


"x=v_0t,"

here, "x=150\ m" is the distance to the target, "v_0=500\ \dfrac{m}{s}" is the initial velocity of the bullet, t is the flight time of the bullet.

Then, from this equation we can calculate "t":


"t=\dfrac{x}{v_0}=\dfrac{150\ m}{500\ \dfrac{m}{s}}=0.3\ s."

Finally, from the another kinematic equation we can find the bullet drop:


"h=\dfrac{1}{2}gt^2=\dfrac{1}{2}\cdot 9.8\ \dfrac{m}{s^2}\cdot (0.3\ s)^2=0.441\ m."

Answer:

"h=0.441\ m."